OpenStudy (i<3mybaby):

The volume of a cylinder is 289 cm2 and the height is 21 cm. Find the diameter. Round the answer to the nearest whole number. A. 2 cm B. 3 cm C. 6 cm D. 4 cm

OpenStudy (anonymous):

a cylinder's volume is \[\pi r^2 * h = v\] so that means that \[\pi r^2 * 21 = 289\] \[\pi r^2= 12.0416\] \[r^2 = 4\] and the root of 4 is 2, so the correct answer is A, 2cm. When you dont know how to solve volume problems simply use what information you have, put it in the volume equation and reverse engineer yourself an answer